OpenStudy (anonymous):

d/dx (tanx-1)/(secx)

OpenStudy (dumbcow):

\[\frac{\tan x -1}{\sec x} = \frac{\tan x}{\sec x} - \frac{1}{\sec x} = \sin x -\cos x\] \[(\sin x - \cos x)' = \cos x +\sin x\]